# Evolving from Single-modal to Multi-modal Facial Deepfake Detection: A Survey

### v1

@@ Abstract
This survey addresses the critical challenge of deepfake detection amidst the rapid advancements in artificial intelligence. As AI-generated media, including video, audio and text, become more realistic, the risk of misuse to spread misinformation and commit identity fraud increases. Focused on face-centric deepfakes, this work traces the evolution from traditional single-modality methods to sophisticated multi-modal approaches that handle audio-visual and text-visual scenarios. We provide comprehensive taxonomies of detection techniques, discuss the evolution of generative methods from auto-encoders and GANs to diffusion models, and categorize these technologies by their unique attributes. To our knowledge, this is the first survey of its kind. We also explore the challenges of adapting detection methods to new generative models and enhancing the reliability and robustness of deepfake detectors, proposing directions for future research. This survey offers a detailed roadmap for researchers, supporting the development of technologies to counter the deceptive use of AI in media creation, particularly facial forgery. A curated list of all related papers can be found at https://github.com/qiqitao77/Awesome-Comprehensive-Deepfake-Detection.
@@ 1 Introduction
Deepfake, encompassing video, audio, and text, has been transformed through advanced artificial intelligence technologies such as Variational Auto-Encoders (VAEs) [1], Generative Adversarial Networks (GANs) [2], and Diffusion Models (DMs) [3] to achieve unprecedented realism. Given that face-related deepfakes may have more severe societal and ethical ramifications than synthetic media of non-facial content, this work specifically focuses on facial deepfakes. It covers various facial manipulations such as face swapping, attribute editing, and face reenactment [4]. Those manipulations not only enhance personalized digital interactions on social platforms and mobile apps but also drive innovations in the entertainment industry, such as digitally resurrecting historical figures [^1] https://edition.cnn.com/2024/02/12/asia/suharto-deepfake-ai-scam-indonesia-election-hnk-intl/index.html or creating virtual actors  [^2] https://edition.cnn.com/2021/08/06/tech/tom-cruise-deepfake-tiktok-company/index.html.
Recent years have witnessed significant advances in deepfake generation, particularly with the emergence of diffusion models, which have revolutionized the creation of synthetic content. This evolution has greatly obscured the distinction between real and fake media, complicating the task for human observers and deepfake detectors. Furthermore, the scope of deepfake technology has expanded beyond single-modality (video or audio alone) to intricate multi-modal manipulations[5][6][7]. These manipulations blend fake audio, video, and text, introducing additional complexity to the detection process. Correspondingly, the deepfake detection mechanisms must evolve to address not only more sophisticated generation techniques but also the challenges posed by the integration of multiple media types.
Progress in deepfake detection over the past two years has addressed the challenges posed by these advanced generative methods and diverse modalities. Researchers have responded to the enhanced realism of synthetic content by innovating detection strategies at different stages of the detection pipeline, including the input level, feature level, and learning level. Some advances leverage new deep learning models, such as transformers [8] and vision-language models [9]. To address multi-modal deepfake detection, various fusion techniques, such as [10, 11, 12, 13], have been developed to utilize knowledge from different modalities effectively. These enhancements refine detection techniques to keep pace with the rapidly evolving quality and diversity of deepfakes, expanding efforts from traditional single-modal to comprehensive multi-modal detection strategies, and from passive to proactive solutions[14][15].
@@
Work Focused Tasks Generative Model Face-centric Multi-modal Released Date
Juefei-Xu et al.[16] Deepfake Generation and Detection GAN & VAE X Mar 2022
Pei et al.[4] Deepfake Generation and Detection GAN & VAE & DM X Mar 2024
Lin et al.[17] AIGC Detection LLM & DM X Feb 2024
Wang et al.[18] Deepfake Detection GAN X Nov 2023
Wang et al.[19] Reliability of Deepfake Detection GAN & VAE X Nov 2022
Ours Deepfake Detection GAN & VAE & DM Jun 2024
@@
The rapid advancements and variations across different works in the field have motivated us to provide the most up-to-date survey, presented in this submission. As demonstrated in Table I, our survey is the first comprehensive review to specifically address facial forgery detection, ranging from single-modal to multi-modal methods. This unique focus distinguishes our survey from previous works in the field. For instance, while [17] explores detection across general fake samples, our survey specifically targets face-related generated media due to its significant potential for societal harm. In contrast to surveys such as [16][4], which allocate much of their discussion to deepfake generation, our work concentrates on detection techniques.
Unlike previous surveys that focused on either facial forgery [16, 4, 18, 19] or multi-modal deepfake [17, 20] separately, our work merges these two crucial areas, offering a holistic view of the current state and challenges of the field.
Our survey is exceptionally current, including the latest research up to June 2024, whereas most prior surveys only cover works released before 2023 [16, 18, 19, 20]. This ensures that we encompass the new emerging detection tasks besides conventional ones. Specifically, we include multi-modal deepfake which represents the trend in real-world scenarios, while existing works[16, 4, 18] merely discuss visually manipulated deepfakes. Besides, we also include the detection of deepfakes by novel generative techniques that are not considered in many previous reviews[16, 4, 18, 19], i.e., synthetic images generated by diffusion models. This survey incorporates the most advanced and state-of-the-art techniques in detection, providing readers with insights into cutting-edge methods and their effectiveness against evolving threats. As such, it serves as an invaluable resource for researchers and practitioners aiming to remain at the forefront of technological advancements in this rapidly evolving field.
@@
More importantly, as illustrated in Figure 3, we offer a more detailed and systematic categorization of the literature than previously available, unlike previous studies  [16] and [4] that categorize deepfake detection methods into less than four types. We contend that our granular examination helps illustrate how different methods enhance detection capabilities in different aspects. By sorting the works according to the manipulation modalities they address, from single-modal to multi-modal detection, our survey clearly differentiates between the techniques used in various contexts.
Within the single-modal category, we distinguish between passive and proactive detection methods, categorizing them based on their objective—whether detecting deepfakes post-generation or preemptively preventing their creation. In passive detection, we recognize the generalization as a core issue, thus dividing the literature into basic and generalized detection. For each type, we further categorize works based on the type of knowledge utilized—artifacts or consistencies—and the stage of application, whether at the input level, model level, or learning level. We also delve into specialized methods designed for newer generative models like DMs, which have not been extensively covered in previous reviews. For proactive detection, we categorize the works based on the generation models they target, such as GANs or DMs.
As the first work comprehensively investigating the multi-modal facial forgery detection domain, predominantly focusing on audio-visual interactions, our survey explores various multi-modal knowledge utilization strategies, including independent learning, joint learning, matching-based approaches, etc.
@@
In summary, our survey not only brings the academic community up to date with the latest developments but also introduces a more nuanced categorization and analysis of facial forgery detection techniques. This structured approach not only aids in understanding the complexities of the field but also inspires new research directions and methodological innovations, making it an essential resource for anyone involved in the study of this direction.
@@ 2.1 Task Formulation
Along with binary classification, spatial and temporal localization are proposed to facilitate detection. Spatial localization applies to image and video input, predicting a mask of the skeptically manipulated region, while temporal localization pertains exclusively to video inputs, requiring the output of a temporal mask indicating the segment(s) manipulated in the given video.
(b) Multi-modal visual-text deepfake dataset $DGM4$, figure from [6].
Figure 1: Multi-modal deepfake detection demonstration.
Multi-modal Audio-Visual Detection[10]. This task tackles the problem when the visual input is video and accompanied by audio stream input, either or both of which can be manipulated. In this case, we denote the paired audio-visual dataset in Figure 1(a) as $\mathcal{D}=\{(\mathbf{X_{i}^{v}},\mathbf{X_{i}^{a}},Y)\}_{i=1}^{N}$, where $\mathbf{X_{i}^{v}}$ is the video defined previously, $\mathbf{X_{i}^{a}}\in\mathbb{R}^{M\times c}$ represents the auditory waveform with $c$ channels and $M$ sample points, and $Y\in\{real,fake\}$ is the binary target label. Similar to the single-modal scenario, multi-modal audio-visual deepfake detection is to obtain an optimal detector $F_{\theta}:\mathbb{R}^{T\times 3\times H\times W}\times\mathbb{R}^{M\times c}\mapsto\{real,fake\}$ incorporating both visual and auditory inputs:
@@ 2.2.1 Single-modal (visual) Deepfake Datasets
Initially, computer graphics techniques are favored to build up deepfake datasets. However, with the advancement of generative AI technologies, GANs, AEs, and DMs have increasingly become preferred methods for deepfake data generation [21, 25, 26, 24, 27, 28, 33, 32]. This shift towards generative AI techniques reflects the progress in the field, offering more sophisticated and convincing deepfakes by leveraging the power of AIGC methods to manipulate or generate content. We show some examples of visual-modal deepfake in Figure 2.
@@
GANs/VAEs based Generated Datasets. The landscape of deepfake detection has been significantly enriched by datasets leveraging GANs and AEs. Pioneering datasets such as FaceForensics++ (FF++)[21], DeepFake Detection Challenge (DFDC) [24], Celeb-DF (and its successor Celeb-DF v2)[25], DeeperForensics-1.0[26], WildDeepfake[27], the Korean Deep-Fake Detection Dataset (KoDF)[28], FFIW10K[29], and ForgeryNet[30] stand at the forefront of this effort. These datasets utilize GANs or VAEs to create synthetic face images that spans various manipulations and scenarios.
@@
FaceForensics++ (FF++)[21] sets a benchmark with its comprehensive collection of manipulations, including face swapping and reenactment through methods like DeepFakes, Face2Face, FaceSwap, and NeuralTextures. A subsequently proposed face swapping technique FaceShifter[23] is applied to the real videos in FF++[21], resulting in a new set of fake samples. DFD[22] enriches deepfake dataset by combining the manipulated and pristine videos recording paid and consenting actors. DFDC [24] challenges detection algorithms with a diverse set of forgery techniques, including eight different methods such as FSGAN [36] and StyleGAN[37], aimed at generating various facial forgeries.
@@
Celeb-DF series [25] harnesses VAEs to craft deepfake images by replicating facial expressions across donor and target faces, pushing the envelope in visual manipulation realism. DeeperForensics-1.0[26] introduces the DeepFake Variational Auto-Encoder (DF-VAE), a learning-based face swapping framework to generate convincing forgeries. KoDF[28] diversifies the dataset pool by incorporating multiple synthesis models and focusing on Korean subjects, addressing the underrepresentation of Asian demographics in prior deepfake detection datasets.
@@
FFIW10K[29] is a large-scale dataset specifically designed for multi-person face forgery detection, highlighting the complexity of detecting alterations among multiple genuine faces. DF-Platter[31] improves deepfake creation techniques using low-resolution videos and emphasizes a gender-balanced, ethnically diverse subject pool. WildDeepfake[27] collects samples from the internet, offering a real-world testbed for detection algorithms. To expand the detection challenge to include spatial and temporal localization tasks, ForgeryNet [30] introduces a comprehensive resource with more than $5,400$ subjects and $15$ manipulation techniques.
During the Covid-19 pandemic, Ko et al. [38] develop a new deepfake dataset featuring face masks by applying masks to images from well-known deepfake collections and noticed a significant drop in detection accuracy of current deepfake detection models when analyzing faces with masks. All these GANs/VAEs based deepfake datasets contribute uniquely to the realm of detecting forgery images, providing researchers with fundamental resources to address the challenges of deepfake detection.
@@
DeepFakeFace[32] employs three distinct DMs—Stable Diffusion v1.5, Stable Diffusion Inpainting, and InsightFace—each generating $30,000$ synthetic images to create a richly diverse dataset tailored for deepfake detection. DiffusionFace[33] stands as the inaugural dataset dedicated to diffusion-based facial forgery, covering a wide range of manipulations from unconditional generation to sophisticated face swapping techniques. It utilizes an impressive array of eleven diffusion models to achieve a new level of realism in forged facial imagery, showcasing the adaptability of diffusion technologies and establishing a benchmark for DMs-based deepfake detection. DiffusionDB-Face and JourneyDB-Face[39] introduce cutting-edge benchmarks for deepfake detection, emphasizing a variety of head poses, facial attributes, styles, and appearances.
The newly developed DiFF[34] is a substantial collection consisting of more than $500,000$ images synthesized using $13$ different generation methods. It utilizes $30,000$ carefully selected textual and visual prompts to ensure that the synthesized images maintain both high fidelity and semantic consistency.
While these datasets primarily support the detection of facial forgeries, the evolving landscape of deepfake content also necessitates broader detection capabilities as seen in datasets like RWDF-23 [40] and Zoom-DF [41]. The RWDF-23 dataset [40], launched in 2023, includes $2,000$ deepfake videos from four online media-sharing platforms, collected between January and May 2023. It encompasses videos in four languages, providing a diverse set of real-world deepfake content. Additionally, this dataset features viewer comments and interactions, and the authors provide a thorough analysis of deepfakes, covering aspects such as creators, manipulation strategies, intentions, and the context of real-world content production. Furthermore, the study in [41] introduces Zoom-DF tailored for remote meeting and video conferencing scenarios. Unlike traditional datasets that focus primarily on facial manipulation, Zoom-DF uniquely concentrates on altering participant movements, expanding the scope of deepfake detection challenges.
Refer to captionRefer to caption
Figure 2: Examples of single-modal deepfake datasets. Images in the top row are fake samples from FF++[21]. Images in the bottom row are generated by Stable Diffusion v1.5. Figure from [32].

### v3

@@ Abstract
This survey addresses the critical challenge of facial deepfake detection amidst the rapid advancements in artificial intelligence. As AI-generated media, including video, audio, and text, become more realistic, the risk of misuse to spread misinformation and commit identity fraud increases. Focused on face-centric deepfakes, this work traces the evolution from traditional single-modality methods to sophisticated multi-modal approaches that handle audio-visual and text-visual scenarios. We provide comprehensive taxonomies of detection techniques, discuss the evolution of generative methods from auto-encoders and GANs to diffusion models, and categorize these technologies by their unique attributes. To our knowledge, this is the first survey of its kind. We also explore the challenges of adapting detection methods to new generative models and enhancing the reliability and robustness of deepfake detectors, proposing directions for future research. This survey offers a detailed roadmap for researchers, supporting the development of technologies to counter the deceptive use of AI in media creation, particularly facial forgery. A curated list of all related papers can be found at https://github.com/qiqitao77/Awesome-Comprehensive-Deepfake-Detection.
@@ 1 Introduction
Deepfake, encompassing video, audio, and text, has been revolutionized by advanced artificial intelligence technologies such as Variational Auto-Encoders (VAEs) [1], Generative Adversarial Networks (GANs) [2], and Diffusion Models (DMs) [3], achieving unprecedented levels of realism. Given that face-related deepfakes may have more severe societal and ethical ramifications than synthetic media of non-facial content, this work specifically focuses on facial deepfakes. Facial deepfakes encompass various facial manipulations such as face swapping, attribute editing, and face reenactment [4]. While these manipulations can enhance personalized digital interactions on social platforms and drive innovations in the entertainment industry, such as digitally resurrecting historical figures [^1] https://edition.cnn.com/2024/02/12/asia/suharto-deepfake-ai-scam-indonesia-election-hnk-intl/index.html or creating virtual actors [^2] https://edition.cnn.com/2021/08/06/tech/tom-cruise-deepfake-tiktok-company/index.html, they also pose significant risks by potentially being used for malicious purposes [^3] https://edition.cnn.com/2024/02/04/asia/deepfake-cfo-scam-hong-kong-intl-hnk/index.html.
Recent years have witnessed significant advances in deepfake generation, including advanced generative methods and various modality manipulations, which is shown in Figure 1. The emergence of DMs, which have revolutionized the creation of synthetic content, has greatly obscured the distinction between real and fake media, complicating the task for human observers and deepfake detectors. Furthermore, the scope of deepfake technology has expanded beyond single-modality (video or audio alone) to intricate multi-modal manipulations[5][6][7]. The blending of fake audio, video, and text adds significant complexity to the detection process, increasing the challenges of identifying these manipulations effectively.
Figure 1: Advances in facial deepfake generation and corresponding detection in recent years.
To address these challenges, corresponding deepfake detection methods have been proposed. Combating with the emergent novel manipulations, researchers have developed various strategies and leveraged advanced larger deep learning models, such as transformers [8] and vision-language models [9], to enhance detection capabilities. To effectively utilize knowledge from different modalities, various fusion techniques [10, 11, 12, 13] have been developed, refining detection methods to keep pace with the rapidly evolving quality and diversity of deepfakes. Furthermore, detection efforts have expanded from passive to proactive solutions [14][15], highlighting the ongoing need to adapt and improve detection methods to counter increasingly sophisticated deepfake techniques.
@@
Work Focused Tasks Face-centric Multi-modal Generative Model Released Date
Juefei-Xu et al.[16] Deepfake Generation and Detection X GAN & VAE Mar 2022
Pei et al.[4] Deepfake Generation and Detection X GAN & VAE & DM Mar 2024
Lin et al.[17] AIGC Detection X LLM & DM Feb 2024
Wang et al.[18] Deepfake Detection X GAN Nov 2023
Wang et al.[19] Reliability of Deepfake Detection X GAN & VAE Nov 2022
Ours Facial Deepfake Detection GAN & VAE & DM Jul 2024
@@
The rapid advancements and variations across different works in the field have motivated us to provide the most up-to-date survey presented in this submission. As demonstrated in Table I, our survey is the first comprehensive review to specifically address facial forgery detection, ranging from single-modal to multi-modal methods. This unique focus distinguishes our survey from previous works in the field. Specifically, this survey possesses following characteristics:
1) It particularly concentrates on facial deepfake detection. In contrast to earlier surveys such as [16][4], which allocate a large part of their discussion to deepfake generation, our work focuses on detection techniques due to the scarcity of comprehensive review for a myriad of existing detection methods. While [17] explores detection across general fake samples, our survey specifically targets face-related generated media due to its significant potential for societal harm.
2) It covers a wider range of facial deepfakes spanning diverse modalities. Specifically, we include multi-modal deepfake, which represents the trend in real-world scenarios, while existing works[16, 4, 18] about facial deepfake merely discuss visually manipulated deepfakes. This extended coverage allows us to provide a more comprehensive understanding of the advancements and issues in deepfake detection.
3) It is exceptionally current, including the latest research up to July 2024, whereas most prior surveys only cover works released before 2023 [16, 18, 19, 20]. This ensures we encompass the detection of fake content resulting from the novel generative techniques, which are not considered in many previous reviews[16, 4, 18, 19], such as those generated by DMs.
@@
4) It offers a more detailed and systematic taxonomy of existing state-of-the-art facial deepfake detection techniques than previously available, as illustrated in Figure 3. Unlike studies such as [16] and [4], which categorize deepfake detection methods into fewer than four types, our granular examination illustrates how different methods enhance detection capabilities in various aspects.
@@
In summary, our survey not only brings the community up-to-date with the latest developments, but also introduces a more nuanced categorization and analysis of facial forgery detection techniques. This structured approach not only aids in understanding the complexities of the field but also inspires new research directions and methodological innovations, making it an essential resource for anyone involved in the study of this direction.
@@ 2.1 Task Formulation
Along with binary classification, spatial and temporal localization are proposed to facilitate detection. Spatial localization applies to either image or video input, predicting a mask of the skeptically manipulated region, while temporal localization pertains exclusively to video inputs, requiring the output of a temporal mask indicating the segment(s) manipulated in the given video.
Multi-modal Audio-Visual Detection[10]. This task tackles the problem when the visual input is video and is accompanied by the audio stream input, either or both of which can be manipulated. In this case, we denote the paired audio-visual dataset in Figure 1(a) as $\mathcal{D}=\{(\mathbf{X_{i}^{v}},\mathbf{X_{i}^{a}},Y)\}_{i=1}^{N}$, where $\mathbf{X_{i}^{v}}$ is the video defined previously, $\mathbf{X_{i}^{a}}\in\mathbb{R}^{M\times c}$ represents the auditory waveform with $c$ channels and $M$ sample points, and $Y\in\{real,fake\}$ is the binary target label. Similar to the single-modal scenario, multi-modal audio-visual deepfake detection is to obtain an optimal detector $F_{\theta}:\mathbb{R}^{T\times 3\times H\times W}\times\mathbb{R}^{M\times c}\mapsto\{real,fake\}$ that incorporates both visual and auditory inputs:
@@ 2.2.1 Single-modal (visual) Deepfake Datasets
Initially, computer graphics techniques are favored to build up deepfake datasets. However, with the advancement of generative AI technologies, GAN, VAE, and DM have become increasingly preferred methods for deepfake data generation [21, 25, 26, 24, 27, 28, 33, 32]. This shift towards generative AI techniques reflects the progress in the field, offering more sophisticated and convincing deepfakes by leveraging the power of AIGC methods to manipulate or generate content.
@@
GANs/VAEs based Generated Datasets. The landscape of deepfake detection has been significantly enriched by datasets leveraging GANs and VAEs. Pioneering datasets such as FaceForensics++ (FF++)[21], DeepFake Detection Challenge (DFDC) [24], Celeb-DF (and its successor Celeb-DF v2)[25], DeeperForensics-1.0[26], WildDeepfake[27], the Korean Deep-Fake Detection Dataset (KoDF)[28], FFIW10K[29], and ForgeryNet[30] stand at the forefront of this effort. These datasets utilize GANs or VAEs to create synthetic face images that span various manipulations.
@@
FaceForensics++ (FF++)[21] sets a benchmark with its comprehensive collection of manipulations, including face swapping and reenactment through methods like DeepFakes, Face2Face, FaceSwap, and NeuralTextures. A subsequently proposed face swapping technique FaceShifter[23] is applied to the real videos in FF++[21], resulting in a new set of fake samples. DFD[22] enriches the deepfake dataset by combining manipulated and pristine videos that record paid and consenting actors. DFDC [24] challenges detection algorithms with a diverse set of forgery techniques, including eight different methods such as FSGAN [37] and StyleGAN[38], aimed at generating various facial forgeries.
@@
Celeb-DF series [25] harnesses VAEs to craft deepfake images by replicating facial expressions across donor and target faces, pushing the envelope in visual manipulation realism. DeeperForensics-1.0[26] introduces the DeepFake Variational Auto-Encoder (DF-VAE), a learning-based face swapping framework to generate convincing forgeries. KoDF[28] diversifies the dataset pool by incorporating multiple synthesis models and focusing on Korean subjects, addressing the underrepresentation of Asian demographics in previous deepfake detection datasets.
@@
FFIW10K[29] is a large-scale dataset specifically designed for multi-person face forgery detection, highlighting the complexity of detecting alterations among multiple genuine faces. DF-Platter[31] improves deepfake creation techniques using low-resolution videos and emphasizes a gender-balanced, ethnically diverse subject pool. WildDeepfake[27] collects samples from the internet, offering a real-world testbed for detection algorithms. To expand the detection challenge to include spatial and temporal localization tasks, ForgeryNet [30] introduces more than $5,400$ subjects and $15$ manipulation techniques. During the Covid-19 pandemic, Ko et al. [39] develop a new deepfake dataset featuring face masks by applying masks to images from well-known deepfake collections and noticed a significant drop in detection accuracy of current deepfake detection models when analyzing faces with masks.
@@
DeepFakeFace[32] employs $3$ distinct DMs—Stable Diffusion v1.5, Stable Diffusion Inpainting, and InsightFace—each generating $30,000$ synthetic images to create a richly diverse dataset tailored for deepfake detection. DiffusionFace[33] is the inaugural dataset dedicated to diffusion-based facial forgery, covering a wide range of manipulations from unconditional generation to sophisticated face swapping techniques. It utilizes $11$ diffusion models to achieve a new level of realism in forged facial imagery, showcasing the adaptability of diffusion technologies and establishing a benchmark for DMs-based deepfake detection. DiffusionDB-Face and JourneyDB-Face[40] introduce cutting-edge benchmarks for deepfake detection, emphasizing a variety of head poses, facial attributes, styles, and appearances. The newly developed DiFF[34] is a substantial collection consisting of more than $500,000$ images synthesized using $13$ different generation methods. It utilizes $30,000$ carefully selected textual and visual prompts to ensure that the synthesized images maintain both high fidelity and semantic consistency.
Recently, Yan et al. create a large-scale and diverse dataset called DF40 [41], featuring $40$ distinct deepfake techniques, including GANs-based and DMs-based methods.. Using DF40, they conduct comprehensive evaluations with 4 standard protocols and $7$ representative detection methods, resulting in more than $2,000$ evaluations. This extensive analysis yields several insightful findings that contribute significantly to the field. For example, CLIP-large emerged as the most powerful baseline, significantly outperforming Xception and demonstrating the ability to generalize to non-face AIGCs even when trained solely on face data.
Although these datasets primarily support the detection of facial forgeries, the evolving landscape of deepfake content also requires broader detection capabilities, as seen in data sets such as RWDF-23 [42], Zoom-DF [43], and AI-Face dataset [44]. The RWDF-23 dataset [42], launched in 2023, includes $2,000$ deepfake videos from four online media-sharing platforms, collected between January and May 2023. It encompasses videos in four languages, providing a diverse set of real-world deepfake content. Additionally, this dataset features viewer comments and interactions, and the authors provide a thorough analysis of deepfakes, covering aspects such as creators, manipulation strategies, intentions, and the context of real-world content production. Furthermore, the study in [43] introduces Zoom-DF designed for remote meetings and video conferencing scenarios. Zoom-DF uniquely concentrates on altering participant movements, expanding the scope of deepfake detection challenges. To assess the fairness of facial deepfake detectors, Li et al. introduce the AI-Face dataset [44]. This groundbreaking dataset is the first of its kind to reach a million-scale and includes demographically annotated AI-generated face images.
(b) Multi-modal visual-text deepfake dataset $DGM4$, figure from [6].
Figure 2: Multi-modal deepfake detection.